OpenStudy (anonymous):

1.determine symetry x^5-31x^3+150x=0

OpenStudy (jdoe0001):

testing for symmetry for y-axis change "x" to "-x" x-axis change "y" to "-y" origin line, change "x" and "y" to "-x" and "-y" if after the change, you end up with the same original equation then it has symmetry to that axis or line for example... let's do the y-axis test "x' to "-x" so \(\large { \underline{x^5-31x^3+150x=f(x)}\qquad \\ \quad \\({\color{brown}{ -x}})^5-31({\color{brown}{ -x}})^3+150({\color{brown}{ -x}})=f({\color{brown}{ -x}}) \\ \quad \\ \begin{cases} (-x)^5\to -x\cdot -x\cdot-x\cdot-x\cdot-x\to &-x^5\\ (-x)^3\to -x\cdot-x\cdot-x\to &-x^3\\ (-x)\to &-x \end{cases}\qquad thus \\ \quad \\ -x^5-31(-x^3)+150(-x)=f(-x)\\ \quad \\\implies \underline{-x^5+31x^3-150x=f(-x)} }\) so.. notice the original equation and f(-x), are not the same... thus, there's no symmetry over the y-axis check for the x-axis and the origin

OpenStudy (anonymous):

so the answer is no symetry?

OpenStudy (anonymous):

@jdoe0001